Westgate Inc. uses a lean manufacturing strategy to manufacture DVR (digital video recorder) players. The company manufacture4. Sale Io 4. Cost
Answer #1
a
Budgeted cell conversion cost per hour 480 per hour =1099200/2290
b
Budgeted cell conversion cost per unit 160 per unit =480*(20/60)
c
1 Raw and in process inventory 62370 =810*77
     Accounts Payable 62370
2 Raw and in process inventory 129600 =810*160
       Conversion costs 129600
3 Finished goods inventory 191970 =810*(77+160)
       Raw and in process inventory 191970
4 Sale Accounts Receivable 322630 =770*419
       Sales 322630
4 Cost Cost of goods sold 182490 =770*(77+160)
        Finished goods inventory 182490